OpenStudy (anonymous):

Gracie Shay wants to buy a new Hummer in 5 years. Gracie estimates the cost of the Hummer will be $28,000. If she invests $12,000 now, at a rate of 6 percent compounded semiannually, she: 1- will have enough money 2- will have exactly $16.000 3- will have $18.000 4- will have $16,126.80 ??????

OpenStudy (gw2011):

The answer is (4): $16,126.80. Here is how you get the answer: A=P(1+i)^n A=The amount P=Principal i=interest rate n=Number of periods the interest applies i=0.06/2=0.03 because the interest is compounded semiannually n=10 because the interest is 3% per period and there are 2 periods per year for 5 years A=(12,000)(1.03)^10 A=(12,000)(1.3439) A=$16,126.80

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

There is a slight difference in doing it differently. the answer for this is 16126.99655. Here is the reason Calculate the effective annual rate (1+0.06/2)^2 -1=0.0609 Now uses the Future value formula 12000(1+0.0609)^5=16126.99655